OpenStudy (anonymous):

Find the set of values of a-"R" , such that tan^2 x + sec^2 x = a holds for some a

OpenStudy (anonymous):

I have solved the question , and i am left with three conditions

OpenStudy (anonymous):

\[\huge a \ge 1 \] or \[\huge a \ge -1\] and \[\huge a \ge \frac{ -5 }{ 4 }\] I am stuck here , wether i should take union or intersection

hartnn (hartnn):

(not verifying whether those conditions are correct)* be careful with the usage of "and"s and "or"s its (1st condition) OR (2nd condition) OR (3rd condition) right ?? because 'a' cannot take 2 values at a time, for example, saying a=10 AND a=11 for one equation is incorrect

OpenStudy (anonymous):

yup , i carefuly used that

hartnn (hartnn):

OR \(\equiv\) Union

OpenStudy (anonymous):

and is intersection?

OpenStudy (anonymous):

so union of 1st and 2nd , let say it is y thus y intersection the third condition

OpenStudy (anonymous):

?

hartnn (hartnn):

yes, AND = Intersection if your "and" is correct, then yes what u said is correct

OpenStudy (anonymous):

yes getting a correct answer and is given the textbook -1 to infinity thank you again ^_^

hartnn (hartnn):

welcome ^_^